# The Mittag-Leffler theorem for proper minimal surfaces and directed meromorphic curves

### v1

@@
{quoting}
[leftmargin=5mm] Abstract  We establish a Mittag-Leffler-type theorem with approximation and interpolation for meromorphic curves $M\to\mathbb{C}^{n}$ ($n\geq 3$) directed by Oka cones in $\mathbb{C}^{n}$ on any open Riemann surface $M$. We derive a result of the same type for proper conformal minimal immersions $M\to\mathbb{R}^{n}$. This includes interpolation in the poles and approximation by embeddings, the latter if $n\geq 5$ in the case of minimal surfaces. As applications, we show that complete minimal ends of finite total curvature in $\mathbb{R}^{5}$ are generically embedded, and characterize those open Riemann surfaces which are the complex structure of a proper minimal surface in $\mathbb{R}^{3}$ of weak finite total curvature.
@@ 1.  Introduction
The last two decades have witnessed the development of the approximation and interpolation theories for minimal surfaces in Euclidean space $\mathbb{R}^{n}$, $n\geq 3$ (see the monograph by Alarcón, Forstnerič, and López [6]). In particular, analogous results to the seminal Runge, Weierstrass, and Mittag-Leffler theorems in complex analysis from the 19th century have been recently established for the family of conformal minimal immersions $M\to\mathbb{R}^{n}$ of any open Riemann surface $M$, leading to a considerable number of applications to the global theory of minimal surfaces. The key tool for this task has been the well known close connection of minimal surfaces with complex analysis via the Weierstrass representation formula, together with the classical theories of approximation and interpolation for holomorphic functions on open Riemann surfaces, as well as the modern Oka theory (see [16] for background). Recall that an immersion $u=(u_{1},\ldots,u_{n}):M\to\mathbb{R}^{n}$ is conformal if and only if the $(1,0)$-differential $\partial u=(\partial u_{1},\ldots,\partial u_{n})$ of $u$ satisfies the nullity condition
@@
Such an immersion $u$ is minimal if and only if it is harmonic if and only if $\partial u$ is holomorphic. Conversely, a nowhere vanishing $\mathbb{C}^{n}$-valued holomorphic 1-form $\phi=(\phi_{1},\ldots,\phi_{n})$ satisfying $\phi_{1}^{2}+\cdots+\phi_{n}^{2}=0$ and the period vanishing condition $\Re\int_{C}\phi=0$ for every closed curve $C$ in $M$ integrates to a conformal minimal immersion $u=\Re\int\phi:M\to\mathbb{R}^{n}$ with $\partial u=\phi$.
@@
The Mittag-Leffler theorem from 1884 [25] states that for any closed discrete subset $E\subset\mathbb{C}$ and any meromorphic function $f$ on a neighbourhood of $E$ there exists a meromorphic function $g$ on $\mathbb{C}$ which is holomorphic on $\mathbb{C}\setminus E$ and satisfies that the difference $g-f$ is holomorphic at every point in $E$. This result was extended by Florack in 1948 [14] (see also Behnke and Stein [10]) to functions on open Riemann surfaces, including interpolation on a closed discrete set; approximation on Runge compact subsets was added by Bishop in 1958 [11]. The natural counterpart of a pole of a meromorphic function in minimal surface theory is a complete end of finite total curvature (see e.g. [26, Chap. 9] or [6, Chap. 4] for background). Indeed, if $u:\overline{\mathbb{D}}_{*}=\{\zeta\in\mathbb{C}:0<|\zeta|\leq 1\}\to\mathbb{R}^{n}$ is a complete conformal minimal immersion of finite total curvature then $\partial u$ is holomorphic on $\overline{\mathbb{D}}_{*}$ and extends meromorphically to $\overline{\mathbb{D}}=\overline{\mathbb{D}}_{*}\cup\{0\}$ with an effective pole at 00. Moreover, $u:\overline{\mathbb{D}}_{*}\to\mathbb{R}^{n}$ is a proper map. Alarcón and López proved in [8] that for any closed discrete subset $E$ of an open Riemann surface $M$ and any complete conformal minimal immersion $u:V\setminus E\to\mathbb{R}^{n}$ on a neighbourhood $V$ of $E$ such that every point in $E$ is an end of finite total curvature of $u$, there is a complete conformal minimal immersion $\tilde{u}:M\setminus E\to\mathbb{R}^{n}$ such that every point in $E$ is an end of finite total curvature of $\tilde{u}$ and the map $\tilde{u}-u$ is continuous (hence harmonic) on $V$ and vanishes at all points in $E$; that is, a Mittag-Leffler theorem for complete immersed minimal surfaces. Their theorem also includes approximation and jet-interpolation. In this paper we shall extend this result in two different directions, namely, we shall establish Mittag-Leffler-type theorems for proper minimal surfaces in $\mathbb{R}^{n}$, embedded if $n\geq 5$ (see Sect. 1.1), and for meromorphic curves in $\mathbb{C}^{n}$ directed by Oka cones (see Sect. 1.2).
1.1.  The Mittag-Leffler theorem for proper minimal surfaces
@@
Let $M$ be an open Riemann surface, $\varnothing\neq E\subset M$ be a closed discrete subset, and $V\subset M$ be a locally connected smoothly bounded closed neighbourhood of $E$ all whose connected components are Runge compact sets. Let $n\geq 3$ be an integer and assume that $u:V\setminus E\to\mathbb{R}^{n}$ is a complete conformal minimal immersion whose restriction to each component of $V\setminus E$ is of finite total curvature; that is, every point in $E$ is a complete end of finite total curvature of $u$. Then there is a proper full conformal minimal immersion $\tilde{u}:M\setminus E\to\mathbb{R}^{n}$ such that $\tilde{u}-u$ is continuous (hence harmonic) at every point in $E$. Moreover, if $n\geq 5$ then we can choose $\tilde{u}$ to be injective (hence a proper embedding).
@@
This result may be upgraded to include Mergelyan approximation on admissible sets, as opposed to mere Runge approximation (cf. Theorem 7.1), and control on the flux as in [8, Theorem 6.1]. Nevertheless, the latter is well understood and we do not repeat it here. The main improvements of Theorem 1.1 with respect to the result by Alarcón and López in [8] are that we can ensure the conformal minimal immersion $\tilde{u}:M\setminus E\to\mathbb{R}^{n}$ to be proper (instead of just complete), and in addition an embedding if $n\geq 5$. Concerning the latter, Theorem 1.1 gives the first general existence result for embedded minimal surfaces in $\mathbb{R}^{5}$ with complete ends of finite total curvature; see (ii) in the following immediate corollary.
@@
Recall that, by a result of Jorge and Meeks [21], a complete minimal end of finite total curvature $u:\overline{\mathbb{D}}_{*}\to\mathbb{R}^{3}$ is embedded (on a neighbourhood of 00) if and only if $\partial u$ has a pole at 00 of order $2$. The special cases of the corollary when $E=\varnothing$ were established in [9] and [5], respectively. Delving into the direction of statement (ii), we record the following corollary to the effect that complete minimal ends of finite total curvature in $\mathbb{R}^{n}$, $n\geq 5$, are generically embedded.
@@
We emphasize that the approximation in the corollary is uniform everywhere on $K\setminus E$. If the map $u|_{\Lambda}$ given in Theorem 1.1 is not injective, then the statement remains to hold except that $\tilde{u}$ cannot be chosen injective. Likewise, if $u:V\setminus E\to\mathbb{R}^{n}$ is not proper, then the statement holds except that $\tilde{u}$ cannot be a proper map. In this case we may choose $\tilde{u}:M\setminus E\to\mathbb{R}^{n}$ to be an almost proper map, hence a complete immersion. [^1] A map $f:X\to Y$ between topological spaces is almost proper if the connected components of $f^{-1}(K)$ are all compact for every compact set $K\subset Y$. If $M$ is a smooth surface, then every almost proper immersion $M\to\mathbb{R}^{n}$ is complete by completeness of $\mathbb{R}^{n}$. Almost proper maps are in some sense the best class of maps $M\to\mathbb{R}^{n}$ that can hit an arbitrary countable set in $\mathbb{R}^{n}$. Thus, we obtain the following result.
@@
By a theorem of Huber [20] (see also Chern-Osserman [13]), if $R$ is a Riemann surface, possibly with nonempty boundary $bR\subset R$, and there is a complete conformal minimal immersion $R\to\mathbb{R}^{n}$ of finite total curvature, then $R$ is of finite conformal type, meaning that $R$ is conformally equivalent to a Riemann surface of the form $\widehat{R}\setminus X$, where $\widehat{R}$ is a compact Riemann surface, possibly with nonempty compact boundary $b\widehat{R}\subset\widehat{R}$, and $X\subset\widehat{R}\setminus b\widehat{R}$ is finite. If $M$ is an open Riemann surface, then a complete conformal minimal immersion $u:M\to\mathbb{R}^{n}$ is of weak finite total curvature, according to López [22, 23], if the restriction $u|_{R}:R\to\mathbb{R}^{n}$ is of finite total curvature for every region $R\subset M$ of finite conformal type. [^2] A connected subset $R\subset M$ of a topological surface $M$ is a region if it is closed and, endowed with the induced topology, a topological surface with possibly nonempty boundary $bR\subset R$. Note that if $M$ is of finite topology, then $u$ is of weak finite total curvature if and only if it is of finite total curvature, and in this case $M$ is a compact Riemann surface with finitely many punctures. By the aforementioned theorem of Huber, if an open Riemann surface $M$ admits a complete conformal minimal immersion $M\to\mathbb{R}^{n}$ of weak finite total curvature, then every region $R\subset M$ with compact boundary and finite topology must be of finite conformal type. Theorem 1.1 implies that the converse statement also holds; this is (iii)$\Rightarrow$(ii) in the following characterization result.
@@
The result is known in the case when $M$ is of finite topology. Indeed, (i)$\Leftrightarrow$(ii)$\Rightarrow$(iii) follows from the classical theorems of Huber [20], Chern-Osserman [13] and Jorge-Meeks [21], while (iii)$\Rightarrow$(ii) is ensured by a recent result of Alarcón and López [8, Theorem 1.2] (see also Pirola [27] for $n=3$).
Now assume that $M$ has infinite topology. The implications (iv)$\Rightarrow$(i) if $n\geq 5$ and (i)$\Rightarrow$(ii) are obvious, while (ii)$\Rightarrow$(iii) is again due to Huber [20]. Let us explain how (iii) implies (iv), if $n\geq 5$, and (i). Assuming (iii), $M$ is biholomorphic to an open Riemann surface of the form $M^{\prime}\setminus E$, where $M^{\prime}$ is still an open Riemann surface (perhaps of finite topology) and $E\subset M^{\prime}$ is a (possibly empty) closed discrete subset, such that every proper region in $M^{\prime}\setminus E$ with compact boundary and finite topology is a finitely punctured compact Riemann surface with compact boundary whose punctures are interior points and belong to $E$ (they are the only ends of the region). Theorem 1.1 applied to $M^{\prime}$ and $E$ then furnishes us with a proper full conformal minimal immersion (embedding if $n\geq 5$) $u:M^{\prime}\setminus E\to\mathbb{R}^{n}$ such that every point in $E$ is a complete end of finite total curvature of $u$. In these conditions, the restriction of the immersion to any proper region of $M^{\prime}\setminus E$ as above is complete and of finite total curvature, hence $u$ is of weak finite total curvature. ∎
It is easily seen that on every open Riemann surface $M^{\prime}$ there is a closed discrete subset $E$ such that $M=M^{\prime}\setminus E$ satisfies condition (iii) in Corollary 1.5 and hence, $M$ is the complex structure of a proper minimal surface in $\mathbb{R}^{3}$ of weak finite total curvature and of a properly embedded minimal surface in $\mathbb{R}^{5}$ of weak finite total curvature. There are several restrictions (even topological ones) on an open Riemann surface $M$ of finite conformal type (thus, conformally equivalent to a finitely punctured compact Riemann surface with empty boundary) to be the complex structure of a complete embedded minimal surface of finite total curvature in $\mathbb{R}^{3}$; see e.g. [19, Sect. 3] or [24, Sect. 5]. It is an open question, likely very difficult, whether every such $M$ admits a complete conformal minimal embedding of finite total curvature into $\mathbb{R}^{4}$ or even into $\mathbb{R}^{5}$.
@@ 1.2.  The Mittag-Leffler theorem for directed meromorphic curves
meaning that the complex derivative $F^{\prime}$ of $F$ with respect to any local holomorphic coordinate on $M$ assumes values in ${\bf A}_{*}={\bf A}\setminus\{0\}$. Equivalently, for any holomorphic 1-form $\theta$ vanishing nowhere on $M$, the range of the holomorphic map $dF/\theta:M\to\mathbb{C}^{n}$ lies in ${\bf A}_{*}$. We refer to [6] for background. More generally, assuming that $A\subset\mathbb{C}^{n}$ is a closed irreducible conical complex subvariety of $\mathbb{C}^{n}$ which is smooth away from 00, a holomorphic immersion $F:M\to\mathbb{C}^{n}$ is directed by $A$, or an $A$-immersion, if its complex derivative $F^{\prime}$ with respect to any local holomorphic coordinate on $M$ takes its values in $A_{*}=A\setminus\{0\}$. Every such subvariety $A$ is algebraic and is the common zero set of finitely many homogeneous holomorphic polynomials by Chow’s theorem, and it turns out that $A_{*}$ is the punctured cone on a connected submanifold $Y$ of $\mathbb{C}\mathbb{P}^{n-1}$. That being so, holomorphic null curves are a special type of directed holomorphic immersions of open Riemann surfaces into $\mathbb{C}^{n}$. Directed holomorphic immersions were studied by Alarcón and Forstnerič [4], who proved an Oka principle with Runge and Mergelyan approximation for holomorphic $A$-immersions, and that every holomorphic $A$-immersion can be approximated uniformly on compact sets by holomorphic $A$-embeddings under the assumption that $A_{*}$ is an Oka manifold, which holds if and only if $Y$ is an Oka manifold (see [4, Proposition 4.5], see [16] for background on Oka theory). Interpolation was added by Alarcón and Castro-Infantes in [3]; we refer to [17, 12, 7] for further recent results in the subject.
@@
Let $A\subset\mathbb{C}^{n}$ $(n\geq 3)$ be a closed irreducible conical subvariety which is not contained in any hyperplane, is smooth away from 00, and such that $A_{*}=A\setminus\{0\}$ is an Oka manifold. Let $M$ be an open Riemann surface, $\varnothing\neq E\subset M$ be a closed discrete subset, and $V\subset M$ be a locally connected smoothly bounded closed neighbourhood of $E$ all whose connected components are Runge compact sets. Let $F:V\setminus E\to\mathbb{C}^{n}$ be a holomorphic $A$-immersion extending meromorphically to $V$ with an effective pole at each point in $E$. Then, for any number $\varepsilon>0$, any closed discrete subset $\Lambda\subset M$ with $\Lambda\subset V\setminus E$, and any map $r:E\cup\Lambda\to\mathbb{N}$ there is a holomorphic $A$-immersion $\widetilde{F}:M\setminus E\to\mathbb{C}^{n}$ satisfying the following conditions.
@@
    If $A$ is as in Theorem 1.6 (vi) and $\varnothing\neq E\subset M$ is closed and discrete, then there is a proper full holomorphic $A$-embedding $M\setminus E\to\mathbb{C}^{n}$ extending meromorphically to $M$ with an effective pole at each point in $E$. (See Proposition 7.4 for a more precise statement.)
@@
    For any closed discrete subset $\varnothing\neq E\subset M$ there is a holomorphic $A$-immersion $F:M\setminus E\to\mathbb{C}^{n}$ extending meromorphically to $M$ with an effective pole at each point in $E$, such that $F(M\setminus E)$ is dense in $\mathbb{C}^{n}$. If in addition $A$ is as in Theorem 1.6 (v), then $F:M\setminus E\to\mathbb{C}^{n}$ can be chosen to be almost proper.
@@
We introduce the notation and the relevant definitions in Section 2. In Section 3, we obtain a semiglobal Mergelyan-type theorem for directed meromorphic immersions with approximation and interpolation (see Theorem 3.1 and Proposition 3.2). This is our main technical result and the key tool to be exploited in the rest of the paper. On the other hand, we prove in Section 4 a desingularization or general position lemma for meromorphic $A$-immersions. By combining these results in a recursive procedure, we establish Theorem 1.6 (the Mittag-Leffler theorem for directed meromorphic immersions) in Sections 5 to 7; see the more precise version in Theorem 7.1. We point out that Theorem 1.1 on minimal surfaces is granted by a very minor modification of the proof of (the somehow more general) Theorem 1.6; one just has to choose the directing cone to be the null quadric in $\mathbb{C}^{n}$ and ignore the imaginary parts of the periods in the construction. This connection between minimal surfaces and directed immersions is already well understood (see e.g. [3, Sect. 7]) and we shall not include the details.

### v2

@@
  • Abstract  We establish a Mittag-Leffler-type theorem with approximation and interpolation for meromorphic curves $M\to\mathbb{C}^{n}$ ($n\geq 3$) directed by Oka cones in $\mathbb{C}^{n}$ on any open Riemann surface $M$. We derive a result of the same type for proper conformal minimal immersions $M\to\mathbb{R}^{n}$. This includes interpolation in the poles and approximation by embeddings, the latter if $n\geq 5$ in the case of minimal surfaces. As applications, we show that complete minimal ends of finite total curvature in $\mathbb{R}^{5}$ are generically embedded, and characterize those open Riemann surfaces which are the complex structure of a proper minimal surface in $\mathbb{R}^{3}$ of weak finite total curvature.
@@ 1. Introduction
The last two decades have witnessed the development of the approximation and interpolation theories for minimal surfaces in Euclidean space $\mathbb{R}^{n}$, $n\geq 3$ (see the monograph by Alarcón, Forstnerič, and López [7]). In particular, analogous results to the seminal Runge, Weierstrass, and Mittag-Leffler theorems in complex analysis from the 19th century have been recently established for the family of conformal minimal immersions $M\to\mathbb{R}^{n}$ of any open Riemann surface $M$. This has led to a considerable number of applications to the global theory of minimal surfaces, a main focus of interest since the works of Osserman in the second half of the 20th century. The key tool for this task has been the well known close connection of minimal surfaces with complex analysis via the Weierstrass representation formula (see e.g. [7, Theorem 2.3.4]), together with the classical theories of approximation and interpolation for holomorphic functions on open Riemann surfaces, as well as the modern Oka theory (see [22, 24] for background). Recall that an immersion $u=(u_{1},\ldots,u_{n}):M\to\mathbb{R}^{n}$ is conformal if and only if the $(1,0)$-differential $\partial u=(\partial u_{1},\ldots,\partial u_{n})$ of $u$ satisfies the nullity condition
@@
Such an immersion $u$ is minimal if and only if it is harmonic if and only if $\partial u$ is holomorphic. Conversely, a nowhere vanishing $\mathbb{C}^{n}$-valued holomorphic 1-form $\phi=(\phi_{1},\ldots,\phi_{n})$ satisfying $\phi_{1}^{2}+\cdots+\phi_{n}^{2}=0$ and the period vanishing condition
$\Re\int_{C}\phi=0\quad\text{for every closed curve $C$ in $M$}$
integrates to a conformal minimal immersion $u=\Re\int\phi:M\to\mathbb{R}^{n}$ with $\partial u=\phi$. We refer a reader who may wish to delve into the background on minimal surface theory to the classical surveys by Osserman [44] and Nitsche [43], as well as to the modern monographs by Colding and Minicozzi [18, 19], Dierkes, Hildebrandt, and Sauvigny [20], and Meeks and Pérez [39, 40], among many others; see [7] for a comprehensive study of minimal surfaces from a complex analytic viewpoint.
@@
The Mittag-Leffler theorem from 1884 [41] states that for any closed discrete subset $E\subset\mathbb{C}$ and any meromorphic function $f$ on a neighbourhood of $E$ there exists a meromorphic function $g$ on $\mathbb{C}$ which is holomorphic on $\mathbb{C}\setminus E$ and satisfies that the difference gfg-f is holomorphic at every point in $E$. This result was extended by Florack in 1948 [21] (see also Behnke and Stein [13]) to functions on open Riemann surfaces, including interpolation on a closed discrete set; approximation on Runge compact subsets was added by Bishop in 1958 [14]. The natural counterpart of a pole of a meromorphic function in minimal surface theory is a complete end of finite total curvature (see e.g. [44, §9] or [7, §4]). Indeed, if $u:\overline{\mathbb{D}}_{*}=\{\zeta\in\mathbb{C}:0<|\zeta|\leq 1\}\to\mathbb{R}^{n}$ is a complete conformal minimal immersion of finite total curvature then $\partial u$ is holomorphic on $\overline{\mathbb{D}}_{*}$ and extends meromorphically to $\overline{\mathbb{D}}=\overline{\mathbb{D}}_{*}\cup\{0\}$ with an effective pole at 0. Moreover, $u:\overline{\mathbb{D}}_{*}\to\mathbb{R}^{n}$ is a proper map. Recall that such a $u$ is of finite total curvature if
$\textrm{TC}(u)=\int_{\mathbb{D}_{*}}KdA>-\infty,$
where $g=u^{*}ds^{2}$ denotes the induced Riemannian metric on $\mathbb{D}_{*}$, $K\leq 0$ is the Gaussian curvature function, and $dA$ is the area element of $(\mathbb{D}_{*},g)$. We refer to [16, 44, 12, 47] for background on the theory of complete minimal surfaces of finite total curvature.
Alarcón and López proved in [9] that for any closed discrete subset $E$ of an open Riemann surface $M$ and any complete conformal minimal immersion $u:V\setminus E\to\mathbb{R}^{n}$ on a neighbourhood $V$ of $E$ such that every point in $E$ is an end of finite total curvature of $u$, there is a complete conformal minimal immersion $\tilde{u}:M\setminus E\to\mathbb{R}^{n}$ such that every point in $E$ is an end of finite total curvature of $\tilde{u}$ and the map $\tilde{u}-u$ is continuous (hence harmonic) on $V$ and vanishes at all points in $E$; that is, a Mittag-Leffler theorem for complete immersed minimal surfaces. Their theorem also includes approximation and jet-interpolation. In this paper we shall extend this result in two different directions, namely, we shall establish Mittag-Leffler-type theorems for proper minimal surfaces in $\mathbb{R}^{n}$, embedded if $n\geq 5$ (see Sec. 1.1), and for meromorphic curves in $\mathbb{C}^{n}$ directed by Oka cones (see Sec. 1.2).
$1.$  The Mittag-Leffler theorem for proper minimal surfaces
@@
Let $M$ be an open Riemann surface, $\varnothing\neq E\subset M$ be a closed discrete subset, and $V\subset M$ be a locally connected smoothly bounded closed neighbourhood of $E$ all whose connected components are Runge compact sets. Let $n\geq 3$ be an integer and assume that $u:V\setminus E\to\mathbb{R}^{n}$ is a complete conformal minimal immersion whose restriction to each component of $V\setminus E$ is of finite total curvature; that is, every point in $E$ is a complete end of finite total curvature of $u$. Then there is a proper full [^1] A conformal minimal immersion $u:M\to\mathbb{R}^{n}$ is full if $f(M)$ is not contained in any hyperplane of $\mathbb{C}^{n}$, where $f=2\partial u/\theta$ for any nowhere vanishing holomorphic 1-form $\theta$ on $M$ [7, Definition 2.5.2]. conformal minimal immersion $\tilde{u}:M\setminus E\to\mathbb{R}^{n}$ such that $\tilde{u}-u$ is continuous (hence harmonic) at every point in $E$. Moreover, if $n\geq 5$ then we can choose $\tilde{u}$ to be injective (hence a proper embedding).
@@
This result may be upgraded to include Mergelyan approximation on admissible sets, as opposed to mere Runge approximation (cf. Theorem 7.1), and control on the flux as in [9, Theorem 6.1]. Nevertheless, the latter is well understood and we do not repeat it here. The main improvements of Theorem 1.1 with respect to the result by Alarcón and López in [9] are that we can ensure the conformal minimal immersion $\tilde{u}:M\setminus E\to\mathbb{R}^{n}$ to be proper (instead of just complete), and in addition an embedding if $n\geq 5$. Note that a generic conformal minimal immersion $M\to\mathbb{R}^{n}$ is complete and nonproper [10], and so finding proper examples is a considerably more difficult task. On the other hand, Theorem 1.1 gives the first general existence result for embedded minimal surfaces in $\mathbb{R}^{5}$ with complete ends of finite total curvature; see (ii) in the following immediate corollary.
@@
Statement (i) contributes to a better understanding of the Sullivan and Schoen-Yau conjectures concerning the conformal properties of proper minimal surfaces in $\mathbb{R}^{3}$; see [46, 42, 11] and also [7, §3.10] for a discussion of these problems. Recall that, by a result of Jorge and Meeks [35], a complete minimal end of finite total curvature $u:\overline{\mathbb{D}}_{*}\to\mathbb{R}^{3}$ is embedded (on a neighbourhood of 0) if and only if $\partial u$ has a pole at 0 of order $2$. The special cases of the corollary when $E=\varnothing$ were established in [11] and [6], respectively. Delving into the direction of statement (ii), we record the following corollary to the effect that complete minimal ends of finite total curvature in $\mathbb{R}^{n}$, $n\geq 5$, are generically embedded.
@@
We emphasize that the approximation in this corollary is uniform everywhere on $K\setminus E$. If the map $u|_{\Lambda}$ given in Theorem 1.1 is not injective, then the statement remains to hold except that $\tilde{u}$ cannot be chosen injective. Likewise, if $u:V\setminus E\to\mathbb{R}^{n}$ is not proper, then the statement holds except that $\tilde{u}$ cannot be a proper map. In this case we may choose $\tilde{u}:M\setminus E\to\mathbb{R}^{n}$ to be an almost proper map, hence a complete immersion. [^2] A map $f:X\to Y$ between topological spaces is almost proper if the connected components of $f^{-1}(K)$ are all compact for every compact set $K\subset Y$. If $M$ is a smooth surface, then every almost proper immersion $M\to\mathbb{R}^{n}$ is complete by completeness of $\mathbb{R}^{n}$. Proper maps are almost proper. Almost proper maps are in some sense the best class of maps $M\to\mathbb{R}^{n}$ that can hit an arbitrary countable set in $\mathbb{R}^{n}$. Thus, we obtain the following result.
@@
By a theorem of Huber [34] (see also Chern-Osserman [16]), if $R$ is a Riemann surface, possibly with nonempty boundary $bR\subset R$, and there is a complete conformal minimal immersion $R\to\mathbb{R}^{n}$ of finite total curvature, then $R$ is of finite conformal type, meaning that $R$ is conformally equivalent to a Riemann surface of the form $\widehat{R}\setminus X$, where $\widehat{R}$ is a compact Riemann surface, possibly with nonempty compact boundary $b\widehat{R}\subset\widehat{R}$, and $X\subset\widehat{R}\setminus b\widehat{R}$ is finite. If $M$ is an open Riemann surface, then a complete conformal minimal immersion $u:M\to\mathbb{R}^{n}$ is of weak finite total curvature, according to López [36, 37], if the restriction $u|_{R}:R\to\mathbb{R}^{n}$ is of finite total curvature for every region $R\subset M$ of finite conformal type. [^3] A connected subset $R\subset M$ of a topological surface $M$ is a region if it is closed and, endowed with the induced topology, a topological surface with possibly nonempty boundary $bR\subset R$. Note that if $M$ is of finite topology, then $u$ is of weak finite total curvature if and only if it is of finite total curvature, and in this case $M$ is a compact Riemann surface with finitely many punctures. By the aforementioned theorem of Huber, if an open Riemann surface $M$ admits a complete conformal minimal immersion $M\to\mathbb{R}^{n}$ of weak finite total curvature, then every region $R\subset M$ with compact boundary and finite topology must be of finite conformal type. Theorem 1.1 implies that the converse statement also holds; this is $\Rightarrow(ii) i$n the following characterization result.
@@
[…]
It is easily seen that on every open Riemann surface $M^{\prime}$ there is a closed discrete subset $E$ such that $M=M^{\prime}\setminus E$ satisfies condition (iii) in Corollary 1.5 and hence, $M$ is the complex structure of a proper minimal surface in $\mathbb{R}^{3}$ of weak finite total curvature and of a properly embedded minimal surface in $\mathbb{R}^{5}$ of weak finite total curvature. There are several restrictions (even topological ones) on an open Riemann surface $M$ of finite conformal type (thus, conformally equivalent to a finitely punctured compact Riemann surface with empty boundary) to be the complex structure of a complete embedded minimal surface of finite total curvature in $\mathbb{R}^{3}$; see e.g. [33, §3] or [38, §5]. It is an open question, likely very difficult, whether every such $M$ admits a complete conformal minimal embedding of finite total curvature into $\mathbb{R}^{4}$ or even into $\mathbb{R}^{5}$.
@@ 1.2.  The Mittag-Leffler theorem for directed meromorphic curves
meaning that the complex derivative $F^{\prime}$ of $F$ with respect to any local holomorphic coordinate on $M$ assumes values in ${\bf A}_{*}={\bf A}\setminus\{0\}$. Equivalently, for any holomorphic 1-form $\theta$ vanishing nowhere on $M$ (such exists by [29, 30]), the range of the holomorphic map $dF/\theta:M\to\mathbb{C}^{n}$ lies in $\bf A$$_{*}$. We refer to [7, §2] for more details. More generally, assuming that $A\subset\mathbb{C}^{n}$ is a closed irreducible conical complex subvariety of $\mathbb{C}^{n}$ which is smooth away from 0, a holomorphic immersion $F:M\to\mathbb{C}^{n}$ is directed by $A$, or an $A$-immersion, if its complex derivative $F^{\prime}$ with respect to any local holomorphic coordinate on $M$ takes its values in $A_{*}=A\setminus\{0\}$. Every such subvariety $A$ is algebraic and is the common zero set of finitely many homogeneous holomorphic polynomials by Chow’s theorem (see e.g. [17]), and it turns out that $A_{*}$ is the punctured cone on a connected submanifold $Y$ of $\mathbb{C}\mathbb{P}^{n-1}$. That being so, holomorphic null curves are a special type of directed holomorphic immersions of open Riemann surfaces into $\mathbb{C}^{n}$. Directed holomorphic immersions were studied by Alarcón and Forstnerič [4], who proved an Oka principle with Runge and Mergelyan approximation for holomorphic $A$-immersions, and that every holomorphic $A$-immersion can be approximated uniformly on compact sets by holomorphic $A$-embeddings under the assumption that $A_{*}$ is an Oka manifold, which holds if and only if $Y$ is an Oka manifold [4, Proposition 4.5]. Interpolation was added by Alarcón and Castro-Infantes in [3]; we refer to [27, 15, 8] for further recent results in the subject. The condition on the subvariety $A_{*}$ to be an Oka manifold arises naturally since holomorphic maps from Stein manifolds to Oka manifolds satisfy the Runge approximation and the Weierstrass interpolation theorems, and more generally, all forms of the Oka principle, in the absence of topological obstructions (see Section 2 for further discussion on the Oka property). These tools shall be extensively used in our proofs. The background on the classical Oka-Grauert theory along with the modern developments in Oka theory are presented in the comprehensive book [24] by Forstnerič, and also in the surveys [28, 25].
@@
Let $A\subset\mathbb{C}^{n}$ $(n\geq 3)$ be a closed irreducible conical subvariety which is not contained in any hyperplane, is smooth away from 0, and such that $A_{*}=A\setminus\{0\}$ is an Oka manifold. Let $M$ be an open Riemann surface, $\varnothing\neq E\subset M$ be a closed discrete subset, and $V\subset M$ be a locally connected smoothly bounded closed neighbourhood of $E$ all whose connected components are Runge compact sets. Let $F:V\setminus E\to\mathbb{C}^{n}$ be a holomorphic $A$-immersion extending meromorphically to $V$ with an effective pole at each point in $E$. Then, for any number $\varepsilon>0$, any closed discrete subset $\Lambda\subset M$ with $\Lambda\subset V\setminus E$, and any map $r:E\cup\Lambda\to\mathbb{N}$ there is a holomorphic $A$-immersion $\widetilde{F}:M\setminus E\to\mathbb{C}^{n}$ satisfying the following conditions.
@@
    If $A$ is as in Theorem 1.6-(vi) and $\varnothing\neq E\subset M$ is closed and discrete, then there is a proper full holomorphic $A$-embedding $M\setminus E\to\mathbb{C}^{n}$ extending meromorphically to $M$ with an effective pole at each point in $E$. (See Proposition 7.4 for a more precise statement.)
@@
    For any closed discrete subset $\varnothing\neq E\subset M$ there is a holomorphic $A$-immersion $F:M\setminus E\to\mathbb{C}^{n}$ extending meromorphically to $M$ with an effective pole at each point in $E$, such that $F(M\setminus E)$ is dense in $\mathbb{C}^{n}$. If in addition $A$ is as in Theorem 1.6-(v), then $F:M\setminus E\to\mathbb{C}^{n}$ can be chosen to be almost proper.
Oka manifolds are special examples of the recently introduced class of Oka-1 manifolds: complex manifolds which are universal targets of holomorphic maps from open Riemann surfaces; see [5, 26] for a precise definition and more information. It remains an open question whether our results (or those in [4]) remain to hold true when the directing cone $A_{*}$ is Oka-1 but not Oka.
@@
We introduce the notation and the relevant definitions in Section 2. In Section 3, we obtain a semiglobal Mergelyan-type theorem for directed meromorphic immersions with approximation and interpolation (see Theorem 3.1 and Proposition 3.2). This is our main technical result and the key tool to be exploited in the rest of the paper. On the other hand, we prove in Section 4 a desingularization or general position lemma for meromorphic $A$-immersions. By combining these results in a recursive procedure, we establish Theorem 1.6 (the Mittag-Leffler theorem for directed meromorphic immersions) in Sections 5 to 7; see the more precise version in Theorem 7.1. We point out that Theorem 1.1 on minimal surfaces is granted by a very minor modification of the proof of (the somehow more general) Theorem 1.6; one just has to choose the directing cone to be the null quadric in $\mathbb{C}^{n}$ and ignore the imaginary parts of the periods in the construction. This connection between minimal surfaces and directed immersions is already well understood (see e.g. [3, §7]) and we shall not include the details.